OpenStudy (anonymous):

27. a. Find the value of each variable to the nearest tenth. b. Find the perimeter of the triangle to the nearest tenth.

OpenStudy (anonymous):

How do I download the attachment for this? I really need help fast. So please help me.

OpenStudy (anonymous):

under the box where you typed: "How do I download the attachment for this? I really need help fast. So please help me." there is an attach file button

OpenStudy (anonymous):

Here is the attachment:

OpenStudy (anonymous):

How can I work this out? Somebody help please.

OpenStudy (anonymous):

one angle is 90 and another is 55 - that means the remaining angle is.....

OpenStudy (anonymous):

these are te options: a = 30, c = 36.6, perimeter = 87.6 a = 36.6, c = 30, perimeter = 87.6 a = 36.6, c = 17.2, perimeter = 74.8 a = 17.2, c = 36.6, perimeter = 74.8

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

bummer they don't even ask for that last angle! sorry phone

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

what do you mean? you cant help me?

OpenStudy (anonymous):

im back - so we know that the 55 angle is equal to tan(55) = opp/adj or tan(55)= a/21 that should give us an answer for side a

OpenStudy (anonymous):

okay? next

OpenStudy (anonymous):

whats the next step?

OpenStudy (anonymous):

well, how do I work this out?: tan(55) = a/21?

OpenStudy (anonymous):

whats c, and the perimeter then?

OpenStudy (anonymous):

\[a ^{2}+b ^{2}=c ^{2}\]

OpenStudy (anonymous):

21^2 + b^2 = c^2 is this right?

OpenStudy (anonymous):

yes and use the answer from tan(55) = a/21 as your b

OpenStudy (anonymous):

tan(55) is just a number - will have to get it from your calculator

OpenStudy (anonymous):

tan(55) = 1.4 right?

OpenStudy (anonymous):

indeed so 1.4 = b/21 multiply both side by 21 to get?

OpenStudy (anonymous):

so, is it 21^2 + 1.4^2 = c^2?

OpenStudy (anonymous):

not yet, 1.4 isnt the length it is the tangent of that angle that we are using to find the length

OpenStudy (anonymous):

umm.... is it 29.4 = b?

OpenStudy (anonymous):

if you used more decimals for tan(55) it is closer to 30 but yes

OpenStudy (anonymous):

so 30^2 + 21^2 = c^2 and you will have all your sides

OpenStudy (anonymous):

900 + 441 = c^2

OpenStudy (anonymous):

good so far

OpenStudy (anonymous):

1,341 = c^2 36.62 = c right?

OpenStudy (anonymous):

that is what I got broganleab is asking almost the same question

OpenStudy (anonymous):

so look at my options that i have listed. Wouldn't the answer be A? The first option.

OpenStudy (anonymous):

or B?

OpenStudy (anonymous):

the hypottinuse hypatonous hippotanoose ...c has to be longer than either a or b

OpenStudy (anonymous):

http://openstudy.com/updates/51659016e4b066fca66105e1 has a better explanation for triangles than mine...if you want to check it out

OpenStudy (anonymous):

Option A or B?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

=) dont forget to award medals ---cause its a competition or something

OpenStudy (anonymous):

okay